OpenStudy (anonymous):

Find the inverse of (3/x) -1 Show steps please

OpenStudy (anonymous):

Medals available here folks!

OpenStudy (akshay_budhkar):

y=3/x-1 y+1=3/x x=3/(y+1) now for inverse interchange x and y therefore your inverse is y=3/(x+1)

OpenStudy (anonymous):

Swap x and y:\[x=\frac{3}{y}-1\]Now solve for y:\[y=\frac{3}{x+1}\]

OpenStudy (anonymous):

Excellent!!